OpenStudy (anonymous):

For Circuit of 10 v & Resistor of 10 ohms,what is the power.

OpenStudy (anonymous):

\[Power = \frac{ Volts^{2} }{ Resistance }\]

OpenStudy (anonymous):

P = I V and I = V/R Therefore: I = 10volts/10ohms = 1Amp so we have P = 1Amp * 10Volts = 10watts

OpenStudy (atabaev):

P=V^2/R =10 watts

OpenStudy (anonymous):

10 Watts